OpenStudy (anonymous):

An object is launched at 19.6 meters per second (m/s) from a 58.8-meter tall platform. The equation for the object's height s at time t seconds after launch is s(t) = –4.9t2 + 19.6t + 58.8, where s is in meters. When does the object strike the ground @zepdrix

zepdrix (zepdrix):

\[\large\rm s(t) = –4.9t^2 + 19.6t + 58.8\]Mmm k so this is the same as the third part of the last problem. We care about the time, t, at which the height of the function, s(t), is zero. So we'll let s(t) be zero and then plug all the goodies into the quadratic formula.\[\large\rm 0 = –4.9t^2 + 19.6t + 58.8\]Plug em in! :O\[\rm t=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):



zepdrix (zepdrix):

What's under the square root? :O It should be 1500 or something